OpenStudy (anonymous):

Answer this and explain = Medal and fan!!!!! Use the function rule to complete the table: -10x + y = 4 x= -2, -1, 0, 1, 2 y=

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

you have to plug in the numbers that x equals first. one at a time. -10*-2 +y-=4

OpenStudy (anonymous):

Ok, Im getting it, but explain a bit more

OpenStudy (anonymous):

that would give you 20+y=4

OpenStudy (anonymous):

then you subtract 20 to both sides and you end up with y=-16

OpenStudy (anonymous):

So I multiply -10 x -2 Or 10 x -2?

OpenStudy (anonymous):

you multiply -10*-2

OpenStudy (anonymous):

Which gives me 20 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And then?

OpenStudy (anonymous):

@MeganChase

OpenStudy (anonymous):

you subtract 20 from 20 which gives you 0 leaving the y by itself and then you subtract 4-20 which gives you -16

OpenStudy (anonymous):

Why do i have to subtract? Is it a rule, that I have to subtract from both sides? I do that every time?

OpenStudy (anonymous):

Well, thanks. i get it more now ;)

OpenStudy (anonymous):

you have to subtract to get y by itself

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

now if you plug it in it should give you 4. with all the equations if you do it right. -10*-2+(-16)=4

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

do you need anymore help?